Grade 11Algebra

the ninth term of an AP is 7. the value of the common difference of the AP. which makes the product a4a9a13 maximum is given by

KAUSHIK SUBRAHMANIAN
10 Years agoGrade 11

1 Answer

Vikas TU
ApprovedApproved Tutor Answer10 Years ago
ninth term of an AP is 7
.i.e
a + 8d = 7 …......................(1)
the product a4a9a13  is:
(a + 3d) (a + 8d) (a + 12d) = M
substitute the value of a from eqn. (1) we get,
(7 – 5d) 7 ( 7 + 4d) = M
dM/dd = 0
7[-5(7 + 4d) + 4 (7-5d)] = 0
40d + 7 = 0
d = -7/40